OpenStudy (anonymous):

The position of a particle moving along the x-axis is given by s(t) = 3t^2 + 6. Use the difference quotient to find the velocity, v(t)=____, and acceleration, a(t)=______. How do I solve this? PLease explain! Thank you:)

zepdrix (zepdrix):

The velocity function is the derivative of position. We take our difference quotient (finding the rate of change of position over time)\[\Large\bf\sf \frac{s(t+\Delta t)-s(t)}{(t+\Delta t)-t}\quad=\quad \frac{s(t+\Delta t)-s(t)}{\Delta t}\]This gives us the average rate of change of position (velocity) over some time period. To find velocity as a function, we want the `instantaneous` rate of change. So we'll let our Delta t approach zero using limits.\[\Large\bf\sf \lim_{\Delta t\to 0} \frac{s(t+\Delta t)-s(t)}{\Delta t}\]This will give us our velocity function.

zepdrix (zepdrix):

So ummmm

zepdrix (zepdrix):

\[\Large\bf\sf s(t)=3t^2+6, \qquad\qquad s(t+\Delta t)=3(t+\Delta t)^2+6\] We just need to plug some stuff in.

OpenStudy (anonymous):

ohh so how do we figure out what to plug in? :/

zepdrix (zepdrix):

Try to get comfortable with the difference quotient :O It's the same formula you used for slope in your basic algebra class except now we're using function notation so it looks a little different. \[\Large\bf\sf \color{royalblue}{s(t)=3t^2+6}, \qquad\qquad \color{orangered}{s(t+\Delta t)=3(t+\Delta t)^2+6}\] We want to plug those two things into our difference quotient below, \[\Large\bf\sf \lim_{\Delta t\to 0} \frac{\color{orangered}{s(t+\Delta t)}-\color{royalblue}{s(t)}}{\Delta t}\]

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{\Delta t\to 0} \frac{\color{orangered}{3(t+\Delta t)^2+6}-\color{royalblue}{(3t^2+6)}}{\Delta t}\]Too much maf? :O Confusing?

zepdrix (zepdrix):

Then from there we simplify.

zepdrix (zepdrix):

What are you thinking food? :D

OpenStudy (anonymous):

haha ohhh is that the same thing as lim h->0 f(a+h) - f(a) / h right?

zepdrix (zepdrix):

yes, the limit definition of the derivative.

OpenStudy (anonymous):

erm would 3(t+∆t)(t+∆t)+6-(3t^2+6)/ ∆t ?

OpenStudy (anonymous):

and we need to do something to get rid of the ∆t in the denominator right?

zepdrix (zepdrix):

We start by `trying to` plug Deltat=0 directly in. We see that we have a problem, so yes we'll need to do some algebra: Expand out the brackets, cancel some stuff out, then try to plug Deltat=0 in again.

OpenStudy (anonymous):

ohh okay so 3(t+t) +6-3t^2 +6 / 0 ?

zepdrix (zepdrix):

Yah, we're dividing by 0, bad bad bad. Since that didn't work, we have to back up, and simplify the expression before we can plug the zero in.

OpenStudy (anonymous):

ohh okay... would this be the simplified version? 3(t+∆t)(t+∆t)+6-(3t^2+6) ---------------------- ∆t

OpenStudy (anonymous):

oh wait... 3(t+∆t)(t+∆t)+6-3(t^2+2) ------------------------ ∆t ??

zepdrix (zepdrix):

You need to multiply out those brackets in the first term.

zepdrix (zepdrix):

Don't factor stuff out, try to expand everything.

zepdrix (zepdrix):

\[\Large\bf\sf (a+b)^2\quad=\quad a^2+2ab+b^2\]\[\Large\bf\sf (t+\Delta t)^2\quad=\quad ?\]

OpenStudy (anonymous):

oh so t^2 + 2t∆t + ∆t^2 ?

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{\Delta t\to 0} \frac{{3(t^2+2t\Delta t+\Delta t^2)+6}-(3t^2+6)}{\Delta t}\]Ok good.

OpenStudy (anonymous):

what happens from here?

zepdrix (zepdrix):

Expaaaand! :O

OpenStudy (anonymous):

3t^2-6t∆t+3∆t^2 + 6 - 3t^2 - 6 ------------------------------ ∆t ?? 3t^2 and 6 cancel out and then you get this? -6t∆t + 3∆t^2 ------------ ∆t ??

zepdrix (zepdrix):

Why is that second term -6t∆t? Shouldn't that be +?

OpenStudy (anonymous):

oh sorry yes, + 6t∆t + 3∆t^2 ------------ ∆t ? better? :)

zepdrix (zepdrix):

Ok looks good! Yes good, you remember to distribute your negative sign.

zepdrix (zepdrix):

Hmm what next? :O Any other cancellations we can make?

OpenStudy (anonymous):

yay!:) um 6t+3∆t ?

zepdrix (zepdrix):

Ok good! We've simplified it down to this, \[\Large\bf\sf \lim_{\Delta t\to0}6t+3 \Delta t\]

OpenStudy (anonymous):

so 6t + 3(0) ?

zepdrix (zepdrix):

Yes, good job, we can try plugging in Deltat=0 now since we've cancelled out some stuff. No longer giving us a problem.

zepdrix (zepdrix):

So we've found that, given our position function:\[\Large\bf\sf s(t)=3t^2+6\]Our velocity function will be:\[\Large\bf\sf v(t)=6t\]

OpenStudy (anonymous):

ohh cool!! :) yay!! so now we have to find a(t) right?

zepdrix (zepdrix):

Yes. a(t) will follow the same process. The acceleration function is the derivative of the velocity function,\[\Large\bf\sf a(t)\quad=\quad \lim_{\Delta t\to 0}\frac{v(t+\Delta t)-v(t)}{\Delta t}\]

OpenStudy (anonymous):

okay so 6t^2 + 6t∆t - 6t --------------- ∆t ??

zepdrix (zepdrix):

Hmm where is the square coming from :o

OpenStudy (anonymous):

6t * t ?

zepdrix (zepdrix):

\[\Large\bf\sf v(\color{royalblue}{t})=6(\color{royalblue}{t}), \qquad\qquad v(\color{royalblue}{t+\Delta t})=6(\color{royalblue}{t+\Delta t})\]

zepdrix (zepdrix):

We plug those in. No square this time.

OpenStudy (anonymous):

ohhh i see now :/ oops!! so 6(t+∆t) - 6(t) ----------- ∆t ?

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

okay and that becomes 6t + 6∆t - 6t / ∆t = 6∆t/∆t = 6 ?

zepdrix (zepdrix):

yay good job.\[\Large\bf\sf a(t)=6\]

OpenStudy (anonymous):

yay!! thank you!!! :D